A horizontal bracket ABC consists of two perpendicular arms AB of a length 0.75 m and BC of a length 0.5 m. The bracket has a solid, circular cross section with a diameter equal to 65 mm. The bracket is inserted in a frictionless sleeve at A (which is slightly larger in diameter), so it is free to rotate about the z0 axis at A and is supported by a pin at C.

Moments are applied at point C M1= 1.5 kN m in the x direction and M2 =1.0 kN m acts in the -z direction. Considering only the moments M1 and M2, calculate the maximum tensile stress, the maximum compressive stress, and the maximum in-plane shear stress at point p, which is located at support A on the side of the bracket at mid-height.

0.75 m Frictionless sleeve embedded in support 20 0.5 m -65 mm Cross section at A
Answer #1

Given data: diameter (d) - 65mm = .065m Length of Arm CAB), b2= 0.75m Length of Arm CBC), by = 0.5m M, = 1.5 km M2 = :0 KN.M0x=0 y=o 0 exy = 0 » Exy = Id -2x (2.289 x104) 8.762 x105 x.065 = -0.8038 mpa ~ -0.804 mpa Zmax = Vox-by) ?t zxył = Vzxy? = (